How do you factor x2 +6x+9
Found 2 solutions by Fombitz, mxgirl22:
Answer by Fombitz(32388)
Answer by mxgirl22(39)



How do you factor x^2+6x+9

**Note: I think you meant x squared. That's why I changed the problem.
------------------------------------------------
1st.) We need to find what will give us 9 when we multiply & will also give us 6 when we add those two numbers together.

I like to look at it like this:

[_] + [_] = 6
[_] * [_] = 9

We know only 1*9=9 and 3*3=9 so from knowing that we can fill in the boxes:

3+3=6
3*3=9


---------------------------------
2nd.) Use the information from above-

x^2+6x+9

x^2+3x+3x+9
(x+3)(x+3)
----------------------------------
3rd.) Check by FOILing (First, Outer, Inner, Last)

(x+3)(x+3)

Distribute left "x" to right "x" and to right "3".

x^2+3x
Distribute left "3" to right "x" and right "3".
3x+9

So it's
x^2+3x+3x+9
Combine like terms.

x^2+6x+9

Gives us our original problem.

---------------------------------------
4th.) Answer.

Factored.
(x+3)(x+3)